raoul has to finish a 473 page book in a week. he decides to read the same number of pages each weekday and 30 extra pages on saturday and on sunday. how many pages will he have to read on wednesday?
Answer by robertb(5830)

Let x = # of pages he has to read on a weekday.
==> x + 30 = # of pages he has to read on a Saturday or a Sunday.
==> 5x + 2(x+30) = 473
==> 5x + 2x + 60 = 473
==> 7x = 413
==> x = 59.
Therefore, he has to read 59 pages on a weekday, particularly on Wednesday.